# Self-Supervised Contrastive Learning for Videos using Differentiable Local Alignment

### v1

@@ 1 Introduction
Action recognition in videos is crucial in computer vision as it enables the analysis and interpretation of human activity. These actions are inherently sequential, consisting of interconnected events. Typically, tasks in video understanding include video classification [1, 2, 3, 4, 5, 6], which involves mapping a global video representation to an action category, or action detection [7, 8, 9, 10, 11], which aims to identify both the spatial and temporal boundaries of specific actions within a video. However, these methods present several challenges: (i) Most approaches require fine-grained annotations, which are time-consuming and labor-intensive; (ii) Most approaches struggle to capture the complex temporal information and causal relationships that vary across diverse videos. To address these challenges, we explore how representation learning based on the pretext-task of video alignment can improve performance in downstream action recognition tasks such as action phase classification and action progression.
@@
Prior self-supervised or weakly-supervised video alignment approaches [12, 13, 14, 15, 16, 17] train on pairs of videos that describe the same action using either cycle-consistency [13], soft dynamic time warping [15, 14], or sequential contrastive learning [16]. We notice that the pretext task of video alignment shares similarities with bioinformatics sequence alignment for studying DNA, RNA, and protein structures. Since some actions are composed of sub-actions or events that occur in a specific order, video alignment shares similarities with protein alignment, as both involve identifying regions of continuity and discontinuity within sequences. There are two main types of alignment algorithms commonly used in bioinformatics: global and local. The Needleman-Wunsch (NW) algorithm [18], a global alignment method, aligns entire sequences from start to finish. This approach is conceptually similar to Dynamic Time Warping (DTW) [19], which aligns sequences of varying lengths by matching their temporal patterns from start to finish. Differentiable versions of NW and DTW algorithms have been established by incorporating smooth maximum/minimum and argmax/argmin functions [20, 21]. In particular, Soft-DTW [21] has been used in prior video alignment approaches [14, 15, 17]. In contrast, the Smith-Waterman (SW) algorithm [22], which implements local alignment, identifies regions of high similarity within sequences and is adaptable to gaps, insertions, and deletions. This capability to handle subsequences makes local alignment particularly appropriate for managing the complexities of action recognition in realistic settings. Recently, bioinformatics research has introduced a differentiable form of the SW algorithm [23, 24], which we adopt for the first time in the context of representation learning to align video sequences.
@@
Specifically, in this work, we introduce a novel loss function termed Local-Alignment Contrastive (LAC) loss, integrated into our proposed end-to-end framework for the video alignment task, as visualized in Fig. 1. Following prior works [16, 17], our framework uses a variation of convolutional and transformer encoders to extract spatio-temporal features from each frame. Our novel LAC loss enables the comparison of video pairs through a consistent and differentiable local alignment loss. This loss specifically accommodates sequences with learned penalties for opening gaps and extending them, combined with a contrastive loss that effectively separates dissimilar frames. We demonstrate that the learned representations of our approach outperform previous methods on various downstream action recognition tasks, such as action phase classification and action phase progression on the Pouring [12] and PennAction [25] datasets.
@@ 2 Related Work
Action Recognition. Over the last decade, researchers have proposed various methods to address video recognition tasks [3, 9, 26, 27, 28]. Video alignment, as a self-supervised pretext task for action recognition, enables the extraction of video representations without requiring labeled data. Previous work has leveraged traditional computer vision techniques to derive these representations. For instance, optical flow [29] computes motion between two frames based on brightness constancy, and the CONDENSATION algorithm [30] employs a probabilistic approach informed by prior motion data. Similarly, Space-Time Interest Points (STIPs) [31] detect salient points in video sequences by adapting image-based point detectors to analyze both spatial and temporal patterns.
@@ 3 Method
This section introduces our novel Local-Alignment Contrastive (LAC) loss function employed within our framework where video alignment serves as the pretext task, as visualized in Fig. 2. Specifically, we will derive the forward and backward passes of the differentiable local alignment loss and explain how we ensure consistency within our method.
@@ 3.1 Notations
Let $V^{1}$ and $V^{2}$ represent a pair of video frames. Following [16], we preprocess the frames via temporal random cropping to generate two cropped videos, each consisting of $T$ frames. We denote these cropped videos as $(V^{1},S^{1})$ and $(V^{2},S^{2})$, where $S\in\mathbb{R}^{T}$ contains the indices of the sampled frames. Next, we apply augmentations, resulting in $\widetilde{V}^{1}$ and $\widetilde{V}^{2}$. Finally, these frames are processed through an encoder to obtain embeddings $Z^{1}$ and $Z^{2}$. Each embedding $Z\in\mathbb{R}^{T\times E}$ represents a sequence with length $T$ and embedding dimension $E$. The objective is to train the encoder so that frames from $V^{1}$ and $V^{2}$ depicting the same action have minimized distances between the corresponding embeddings.
@@ 3.2 Background
Prior approaches [14, 15] use DTW as the principal method for temporal alignment loss. Given a cost matrix $d_{i,j}$ that measures the cost of matching elements between embeddings $Z^{1}$ and $Z^{2}$, DTW seeks the optimal path that minimizes this cost. The feasible paths must satisfy constraints of matching endpoints, monotonicity, and continuity. Initially, the computational complexity of DTW grows exponentially with the length of the sequences due to the combinatorial explosion of possible alignment paths. However, by using dynamic programming, DTW’s complexity reduced to polynomial time. The DTW algorithm can be formulated as follows:
$D_{i,j}=d_{i,j}+\min\left\{\begin{array}[]{l}D_{i-1,j-1}\\ D_{i-1,j}\\ D_{i,j-1}\\ \end{array}\right.$ (1)
where $D_{i,j}$ represents the cumulative distance up to the point $(i,j)$ in the cost matrix, and $d_{i,j}$ denotes the local distance between elements $Z^{1}_{i}$ and $Z^{2}_{j}$. The non-differentiable nature of the $\min$ operator in the DTW formulation presents challenges for gradient-based optimization techniques. To address this, prior methods [14, 15] utilize Soft-DTW [21], which replaces the discrete $\min$ operator with a differentiable/smoothed version denoted as $\min^{\gamma}$. As the parameter $\gamma$ approaches zero, he behavior of $\min^{\gamma}$ approximates that of the discrete $\min$ operator. This substitution enables the use of gradient-based optimization in DTW while preserving its alignment capabilities.
@@ 3.3 Differentiable Local-Alignment
Despite the advantages offered by Soft-DTW [21], it inherently maintains the global alignment characteristic, ensuring continuity throughout the sequence. This attribute of global alignment may prove suboptimal in scenarios where localized alignment strategies, such as those implemented by the SW algorithm [22], are preferable. The SW algorithm [22] supports localized matching, providing enhanced flexibility and increased sensitivity to similarities within subsequences. In response, we propose a differentiable SW local-alignment approach as our method for temporal alignment loss.
We consider a pair of embeddings, $(Z^{1},S^{1})$ and $(Z^{2},S^{2})$, with corresponding sampled indices, tasked with determining the optimal alignment between these embeddings. We begin by calculating a similarity matrix, $S\in\mathbb{R}^{T\times T}$, using the inverted Euclidean distance to yield higher values that indicate increased similarity between the points in $Z^{1}$ and $Z^{2}$. Utilizing dynamic programming, the SW algorithm effectively determines the optimal alignment by maximizing a score derived from pairwise similarities within the matrix. The alignment algorithm can be formally expressed as follows:
@@ 4.1 Results
We apply the same four metrics to the Pouring dataset [12]. For the PennAction dataset [25], we evaluate each metric across all action categories and report the average results. To ensure a fair comparison, we replicated the evaluations of previous approaches using the GitHub repositories 111github.com/google-research/google-research/tcc, github.com/trquhuytin/LAV-CVPR21,
@@ 4.2 Ablation Study
Hyperparameter of LAC Loss. Table 5 presents the optimal performance results obtained at the smoothing parameter $\gamma=0.8$. Table 5 further demonstrates that integrating local alignment loss into the total loss significantly enhances performance and achieves the highest scores across all metrics compared to using contrastive loss alone.

### v2

@@ 1 Introduction
Video understanding and analysis have garnered significant attention in computer vision research in recent years [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. These tasks require models to capture not only spatial information within individual frames but also the complex temporal dynamics that evolve across sequences of frames. Traditional supervised learning methods have achieved considerable success in advancing performance; models such as those presented in [2, 4, 10] are able to predict action categories effectively. However, these methods present several challenges: (i) Most approaches require fine-grained annotations, which are time-consuming and labor-intensive; (ii) Most approaches struggle to capture the complex temporal information and causal relationships that vary across diverse videos. To address these challenges, we explore how representation learning based on the pretext-task of video alignment can improve performance in downstream action recognition tasks such as phase classification and phase progression.
@@
Prior self-supervised or weakly-supervised video alignment approaches [11, 12, 13, 14, 15, 16] train on pairs of videos that describe the same action using either cycle-consistency [12], soft dynamic time warping [14, 13], or sequential contrastive learning [15]. We notice that the pretext task of video alignment shares similarities with bioinformatics sequence alignment for studying DNA, RNA, and protein structures. Since some actions are composed of sub-actions or events that occur in a specific order, video alignment shares similarities with protein alignment, as both involve identifying regions of continuity and discontinuity within sequences. There are two main types of alignment algorithms commonly used in bioinformatics: global and local. The Needleman-Wunsch (NW) algorithm [17], a global alignment method, aligns entire sequences from start to finish. This approach is conceptually similar to Dynamic Time Warping (DTW) [18], which aligns sequences of varying lengths by matching their temporal patterns from start to finish. Differentiable versions of NW and DTW algorithms have been established by incorporating smooth maximum/minimum and argmax/argmin functions [19, 20]. In particular, Soft-DTW [20] has been used in prior video alignment approaches [13, 14, 16]. In contrast, the Smith-Waterman (SW) algorithm [21] implements local alignment, identifying regions of high similarity within sequences and allowing for gaps, insertions, and deletions. This ability to handle subsequences makes local alignment particularly suitable for modeling the complexity of actions in real-world video data, where actions may not always follow rigid, continuous sequences.
@@
Specifically, in this work, we introduce a novel loss function termed Local-Alignment Contrastive (LAC) loss, integrated into our proposed end-to-end framework for the video alignment task, as visualized in Figure 1. Following prior works [15, 16], our framework uses a variation of convolutional and transformer encoders to extract spatio-temporal features from each frame. However, rather than relying on a differentiable global alignment loss like Soft-DTW [20], our novel LAC loss enables the comparison of video pairs through a consistent and differentiable local alignment loss. This loss specifically accommodates sequences with learned penalties for opening gaps and extending them, combined with a contrastive loss that effectively separates dissimilar frames. We demonstrate that the learned representations of our approach outperform previous methods on various downstream action recognition tasks, such as action phase classification and action phase progression on the Pouring [11] and PennAction [22] datasets. Code and models are available at: https://github.com/keynekassapa13/LAC.
@@ 2 Related Work
Action Recognition. Over the last decade, researchers have proposed various methods to address video recognition tasks [2, 4, 1, 3, 5], including action classification [23, 24, 2, 6, 8, 9], which involves mapping a video to an action category, and action segmentation [25, 26, 4, 27, 28], which aims to identify the spatial and temporal boundaries of specific actions within a video. However, these methods often rely heavily on extensive labeled data, which is time-consuming and labor-intensive to obtain. To address this limitation, self-supervised learning methods have been explored, enabling the extraction of video representations without requiring labeled data. Video alignment, as a self-supervised pretext task for action recognition, leverages temporal consistency to learn meaningful representations. Previous work has utilized traditional computer vision techniques to derive these representations; for instance, optical flow [29] computes motion between frames based on brightness constancy, the CONDENSATION algorithm [30] employs a probabilistic approach informed by prior motion data, and Space-Time Interest Points (STIPs) [31] detect salient points by analyzing both spatial and temporal patterns in video sequences.
@@
Global vs. Local Alignment. Global alignment methods, such as the NW algorithm [17] and DTW [18], aim to align entire sequences from start to finish to achieve the best overall match; differentiable versions have been developed [20, 19] and utilized in prior video alignment approaches [13, 14, 16]. In contrast, the SW algorithm [21] focuses on identifying the most similar subsequences within larger sequences. Recent work by Petti et al. [32] introduced a differentiable SW algorithm in a supervised framework that jointly learns alignments and improves protein structure prediction. [33] shows the power of embedding sequences into high-dimensional spaces within a supervised framework for more precise local alignment. Our approach employs a differentiable version of SW in self-supervised framework to improve model performance on video understanding and alignment tasks compared to state-of-the-art temporal global alignment methods.
@@ 3 Method
This section introduces our novel Local-Alignment Contrastive (LAC) loss function employed within our framework where video alignment serves as the pretext task, as visualized in Fig. 2. Specifically, we derive the forward and backward passes of the differentiable local alignment loss and explain how we ensure consistency within our method.
@@ 3.1 Notations
Let $V^{1}$ and $V^{2}$ represent two videos, each consisting of a sequence of frames. Specifically, $V^{1}=\{v^{1}_{t}\}_{t=1}^{T_{1}}$ and $V^{2}=\{v^{2}_{t}\}_{t=1}^{T_{2}}$, where $v^{1}_{t}$ and $v^{2}_{t}$ denote individual frames at time $t$, and $T_{1}$ and $T_{2}$ are the total number of frames in each video. Following [15], we preprocess the videos using temporal random cropping to generate cropped versions $V_{c}^{1}$ and $V_{c}^{2}$. The indices of the sampled frames during cropping are recorded in $S^{1}$ and $S^{2}$, where $S^{1}\subseteq\{1,\dotsc,T_{1}\}$ and $S^{2}\subseteq\{1,\dotsc,T_{2}\}$. Next, we apply data augmentations to the cropped videos, resulting in $\widetilde{V}^{1}$ and $\widetilde{V}^{2}$. These augmented videos are then processed through an encoder to obtain embeddings $Z^{1}$ and $Z^{2}$. Each embedding $Z^{i}\in\mathbb{R}^{T_{i}\times E}$ represents a sequence of length $T_{i}$ with an embedding dimension $E$. The objective is to train the encoder to minimize the distance between embeddings $Z^{1}$ and $Z^{2}$ corresponding to videos depicting the same action, as determined by our loss.
@@ 3.2 Background
Prior approaches [13, 14] use DTW as the main method for the temporal alignment loss. Given embeddings $Z^{1}$ and $Z^{2}$, we compute a cost matrix $C\in\mathbb{R}^{T_{1}\times T_{2}}$, where each element $C_{i,j}$ measures the local distance between $Z^{1}_{i}$ and $Z^{2}_{j}$. DTW seeks the optimal alignment path through this cost matrix that minimizes the total cumulative cost. The feasible paths must satisfy constraints of matching endpoints, monotonicity, and continuity. While the number of possible alignment paths grows exponentially with the lengths of the sequences, dynamic programming allows DTW to compute the optimal path efficiently in quadratic time $O(T_{1}T_{2})$. The DTW algorithm is formulated recursively as:
$D_{i,j}=C_{i,j}+\min\left\{\begin{array}[]{l}D_{i-1,j-1}\\ D_{i-1,j}\\ D_{i,j-1}\\ \end{array}\right.$ (1)
where $D_{i,j}$ represents the accumulated cost up to point $(i,j)$, and $C_{i,j}$ is the local cost between elements $Z^{1}_{i}$ and $Z^{2}_{j}$. The non-differentiable nature of the DTW formulation presents challenges for gradient-based optimization techniques. To address this, prior methods [13, 14] utilize Soft-DTW [20] which replaces the discrete $\min$ operator with a differentiable, smoothed version denoted as $\text{softmin}^{\gamma}$. As the smoothing parameter $\gamma$ approaches zero, the behavior of $\text{softmin}^{\gamma}$ approximates that of the discrete $\min$ operator. This substitution enables the use of gradient-based optimization in DTW while preserving its alignment capabilities.
@@ 3.3 Differentiable Local-Alignment
Despite the advantages offered by Soft-DTW [20], it inherently maintains the global alignment characteristic, ensuring continuity throughout the entire sequence. This attribute of global alignment may prove suboptimal in scenarios where localized alignment strategies, such as those implemented by the SW algorithm [21], are preferable. The SW algorithm [21] supports localized matching, providing enhanced flexibility and increased sensitivity to similarities within subsequences. Therefore, we propose a differentiable SW local-alignment approach as our method for temporal alignment loss.
We consider a pair of embeddings, $(Z^{1},S^{1})$ and $(Z^{2},S^{2})$, with corresponding sampled indices, tasked with determining the optimal alignment between these embeddings. We begin by calculating a similarity matrix, $S\in\mathbb{R}^{T\times T}$, using the inverted Euclidean distance to yield higher values that indicate increased similarity between the points in $Z^{1}$ and $Z^{2}$. Utilizing dynamic programming, the SW algorithm effectively determines the optimal alignment by maximizing a score derived from pairwise similarities within the matrix. The alignment algorithm is formally expressed as follows:
@@ 4.1 Results
We apply the same four metrics to the Pouring dataset [11]. For the PennAction dataset [22], we evaluate each of the 13 action categories using the same four metrics and report the averaged results across all categories. To ensure a fair comparison (e.g., consistent data splits, identical preprocessing), we replicated the evaluations of previous approaches using the GitHub repositories 111github.com/google-research/google-research/tcc, github.com/trquhuytin/LAV-CVPR21,
@@ 4.2 Ablation Study
Hyperparameter of LAC Loss. Table 5 presents the optimal performance results obtained at the smoothing parameter $\gamma=0.8$. Table 5 demonstrates that adding the local alignment loss ($\mathcal{L}_{l}$) to the contrastive loss ($\mathcal{L}_{c}$), the classification accuracy increases by 1.72, progression accuracy by 2.12, and the Tau score by 2.57.
@@ 5 Conclusion
Supplementary Materials
A  Architecture Details
The complete architecture of our encoder is shown in Table 6. The encoder consists of two main components: the Spatial Encoder and the Temporal Encoder. The Spatial Encoder is responsible for extracting spatial features from individual frames using ResNet-50 [35] with pretrained weights. This extraction process involves a series of convolutional layers, batch normalization, ReLU activations, max pooling, and bottleneck layers. Subsequently, the Temporal Encoder processes these spatial features across multiple frames to capture temporal dependencies. It employs fully connected layers with dropout and ReLU activations to enhance feature learning and prevent overfitting. A positional encoder is integrated to incorporate sequence information, and transformer encoder layers are utilized to model temporal relationships effectively. The final linear layer reduces the dimensionality of the embeddings to 128.
Model Layer Parameters Output Size
Spatial Encoder Conv2d $647x7,$ 64 $112\times 112\times 64$
BatchNorm2d + ReLU -
MaxPool2d $33x$ 3 $56\times 56\times 64$
Bottleneck conv2_x $\left[\begin{array}[]{c}1\times 1,64\\ 3\times 3,64\\ 1\times 1,256\end{array}\right]\times 3$ $56\times 56\times 256$
Bottleneck conv3_x $\left[\begin{array}[]{c}1\times 1,128\\ 3\times 3,128\\ 1\times 1,512\end{array}\right]\times 4$ $28\times 28\times 512$
Bottleneck conv4_x $\left[\begin{array}[]{c}1\times 1,256\\ 3\times 3,256\\ 1\times 1,1024\end{array}\right]\times 3$ $14\times 14\times 1024$
Conv3d $5123x3x3,$ 512 $2\times 10\times 10\times 512$
BatchNorm3d + ReLU -
AdaptiveMaxPool3d 1 $1\times 1\times 1\times 512$
Temporal Encoder Dropout $\text{p}=0.1$ $64\times 512$
Linear 512
ReLU -
Linear 256 $64\times 256$
Positional Encoder - $2\times 32\times 256$
Transformer Encoder $\left[\begin{array}[]{c}256,8,1024\end{array}\right]\times 2$ $2\times 32\times 256$
Linear 128 $64\times 128$
Table 6: Summarized Composition of the Spatial and Temporal Encoder Architectures. The parameter structure is as follows: For convolution layers, the format is kernel size and number of output channels (e.g., $7\times 7,64$). For pooling layers, the format is kernel size (e.g., $3\times 3$). For other layers, specific parameters are listed (e.g., $p=0.1$ for dropout). For the Transformer Encoder, the format is the dimension of input and output vectors, number of attention heads, and dimension of the feedforward network hidden layer (e.g., $\left[\begin{array}[]{c}256,8,1024\end{array}\right]\times 2$).
B  Differentiable Local-Alignment in Detail
We provide a detailed derivation of the backward recursion used in differentiable local alignment algorithms. Following the approach described in [33], we transform the discrete max function into a differentiable form, specifically a smooth max function. This smoothe max function is mathematically represented as follows:
$\max{}^{\gamma}(u_{1},\ldots,u_{n}):=\gamma\,\log\,\sum_{i=1}^{n}\,e^{u_{i}/\gamma},$ (9)
Equation 9 enables differentiation that results in a softmax function, detailed as follows:
$=\exp\left(\frac{u_{i}-\gamma\log\left(\sum_{j=1}^{n}\exp\left(\frac{u_{j}}{\gamma}\right)\right)}{\gamma}\right)$ (11)
$=\exp\left(\frac{u_{i}}{\gamma}-\log\left(\sum_{j=1}^{n}\exp\left(\frac{u_{j}}{\gamma}\right)\right)\right)$ (12)
$=\exp\left(\frac{u_{i}}{\gamma}\right)\cdot\exp\left(-\log\left(\sum_{j=1}^{n}\exp\left(\frac{u_{j}}{\gamma}\right)\right)\right)$ (13)
$=\frac{\exp\left(\frac{u_{i}}{\gamma}\right)}{\sum_{j=1}^{n}\exp\left(\frac{u_{j}}{\gamma}\right)}$ (14)
This transformation makes the function differentiable, allowing the use of gradient-based optimization in alignment algorithms. The Smith-Waterman (SW) local alignment algorithm [21] can be defined using dynamic programming as follows:
$I_{x_{i,j}}=\max\left\{\begin{array}[]{l}D_{i,j-1}-g_{o_{i,j}}\\ I_{x_{i,j-1}}-g_{e_{i,j}}\end{array}\right.,\quad I_{y_{i,j}}=\max\left\{\begin{array}[]{l}D_{i-1,j}-g_{o_{i,j}}\\ I_{x_{i-1,j}}-g_{o_{i,j}}\\ I_{y_{i-1,j}}-g_{e_{i,j}}\end{array}\right.$ (16)
where $D_{i,j}$ represents the score of the optimal alignment ending at positions $i$ and $j$ in the two sequences being aligned. The scores for initiating or extending a gap are handled by $I_{x_{i,j}}$ and $I_{y_{i,j}}$. Additionally, $g_{o_{i,j}}$ and $g_{e_{i,j}}$ denote the penalties for opening and extending gaps. This method enables differentiation of the smoothed maximum score, denoted as $s*=\max^{\gamma}(0,\max^{\gamma}(D))$, with respect to $S$, $g_{o}$, and $g_{e}$, treating these as parameters sensitive to derivatives. The differentiation of the smoothed maximum score with respect to $S$ can be defined as follows:
$\frac{\partial s^{*}}{\partial S_{i,j}}$ $=\left(\frac{\partial s^{*}}{\partial D_{i,j}}\right)\frac{\partial D_{i,j}}{\partial S_{i,j}}$ (17)
Using the inverse relationship defined in Eq. 16, the differentiation of the smoothed maximum score with respect to $g_{o}$ is described as follows:
$\frac{\partial s^{*}}{\partial g_{o_{i,j}}}$ $=\frac{\partial s^{*}}{\partial I_{x_{i,j}}}\,\frac{\partial I_{x_{i,j}}}{\partial g_{o_{i,j}}}+\frac{\partial s^{*}}{\partial I_{y_{i,j}}}\,\frac{\partial I_{y_{i,j}}}{\partial g_{o_{i,j}}}$ (18)
$=\frac{\partial s^{*}}{\partial I_{x_{i,j}}}\,\frac{\partial\max^{\gamma}\left(D_{i,j-1}-g_{o_{i,j}},I_{x_{i,j-1}}-g_{e_{i,j}}\right)}{\partial g_{o_{i,j}}}$ (19)
$\quad+\frac{\partial s^{*}}{\partial I_{y_{i,j}}}\,\frac{\partial\max^{\gamma}\left(D_{i-1,j}-g_{o_{i,j}},I_{x_{i-1,j}}-g_{o_{i,j}},I_{y_{i-1,j}}-g_{e_{i,j}}\right)}{\partial g_{o_{i,j}}}$ (20)
$=\frac{\partial s^{*}}{\partial I_{x_{i,j}}}\left(-\frac{\partial I_{x_{i,j}}}{\partial D_{i,j-1}}\right)+\frac{\partial s^{*}}{\partial I_{y_{i,j}}}\left(-\frac{\partial I_{y_{i,j}}}{\partial D_{i-1,j}}-\frac{\partial I_{y_{i,j}}}{\partial I_{x_{i-1,j}}}\right)$ (21)
Similar to the above, using the inverse relationship from Eq. 16, the differentiation of the smoothed maximum score with respect to $g_{e}$ is defined as follows:
$\frac{\partial s^{*}}{\partial g_{e_{i,j}}}$ $=\frac{\partial s^{*}}{\partial I_{x_{i,j}}}\frac{\partial I_{x_{i,j}}}{\partial g_{e_{i,j}}}+\frac{\partial s^{*}}{\partial I_{y_{i,j}}}\frac{\partial I_{y_{i,j}}}{\partial g_{e_{i,j}}}$ (22)
$=\frac{\partial s^{*}}{\partial I_{x_{i,j}}}\frac{\partial\max^{\gamma}\left(D_{i,j-1}-g_{o_{i,j}},I_{x_{i,j-1}}-g_{e_{i,j}}\right)}{\partial g_{e_{i,j}}}$ (23)
$\quad+\frac{\partial s^{*}}{\partial I_{y_{i,j}}}\frac{\partial\max^{\gamma}\left(D_{i-1,j}-g_{o_{i,j}},I_{x_{i-1,j}}-g_{o_{i,j}},I_{y_{i-1,j}}-g_{e_{i,j}}\right)}{\partial g_{e_{i,j}}}$ (24)
$=\frac{\partial s^{*}}{\partial I_{x_{i,j}}}\left(-\frac{\partial I_{x_{i,j}}}{\partial I_{x_{i,j-1}}}\right)+\frac{\partial s^{*}}{\partial I_{y_{i,j}}}\left(-\frac{\partial I_{y_{i,j}}}{\partial I_{y_{i-1,j}}}\right)$ (25)
Following [33], we consider backtracking the SW algorithm to set the recursive calculation and see how changes in the current state propagate gradients forward to influence future states. We begin by deriving the expression for $\frac{\partial s^{*}}{\partial D_{i,j}}$ from Eq. 17, which can be obtained as follows:
$\frac{\partial s^{*}}{\partial D_{i,j}}$ $=\frac{\partial s^{*}}{\partial D_{i,j}}+\frac{\partial s^{*}}{\partial D_{i+1,j+1}}\frac{\partial D_{i+1,j+1}}{\partial D_{i,j}}+\frac{\partial s^{*}}{\partial I_{x_{i,j+1}}}\frac{\partial I_{x_{i,j+1}}}{\partial D_{i,j}}+\frac{\partial s^{*}}{\partial I_{y_{i+1,j}}}\frac{\partial I_{y_{i+1,j}}}{\partial D_{i,j}}$ (26)
In this equation, $\frac{\partial s^{*}}{\partial D_{i,j}}$ represents the immediate partial derivative of $s^{*}$ with respect to $D_{i,j}$, reflecting its direct impact. The term $\frac{\partial s^{*}}{\partial D_{i+1,j+1}}\frac{\partial D_{i+1,j+1}}{\partial D_{i,j}}$ accounts for how $D_{i,j}$ affects the future state $D_{i+1,j+1}$. Similarly, $\frac{\partial s^{*}}{\partial I_{x_{i,j+1}}}\frac{\partial I_{x_{i,j+1}}}{\partial D_{i,j}}$ reflects the influence on the future state $I_{x_{i,j+1}}$, and $\frac{\partial s^{*}}{\partial I_{y_{i+1,j}}}\frac{\partial I_{y_{i+1,j}}}{\partial D_{i,j}}$ reflects the influence on the future state $I_{y_{i+1,j}}$.
From Equations 21 and 25, we see that the remaining partial derivatives to be calculated are $\frac{\partial s^{*}}{\partial I_{x_{i,j}}}$ and $\frac{\partial s^{*}}{\partial I_{y_{i,j}}}$. These derivatives can be derived recursively as follows:
$\frac{\partial s^{*}}{\partial I_{y_{i,j}}}$ $=\frac{\partial s^{*}}{\partial D_{i+1,j+1}}\,\frac{\partial D_{i+1,j+1}}{\partial I_{y_{i,j}}}+\frac{\partial s^{*}}{\partial I_{y_{i+1,j}}}\,\frac{\partial I_{y_{i+1,j}}}{\partial I_{y_{i,j}}}$ (27)
$\frac{\partial s^{*}}{\partial I_{x_{i,j}}}$ $=\frac{\partial s^{*}}{\partial D_{i+1,j+1}}\,\frac{\partial D_{i+1,j+1}}{\partial I_{x_{i,j}}}+\frac{\partial s^{*}}{\partial I_{x_{i,j+1}}}\,\frac{\partial I_{x_{i,j+1}}}{\partial I_{x_{i,j}}}+\frac{\partial s^{*}}{\partial I_{y_{i+1,j}}}\,\frac{\partial I_{y_{i+1,j}}}{\partial I_{x_{i,j}}}$ (28)
where $\frac{\partial s^{*}}{\partial I_{y_{i,j}}}$ accounts for the influence of $I_{y_{i,j}}$ on the future states $D_{i+1,j+1}$ and $I_{y_{i+1,j}}$. Similarly, $\frac{\partial s^{*}}{\partial I_{x_{i,j}}}$ reflects the influence of $I_{x_{i,j}}$ on the future states $D_{i,j+1}$ and $I_{x_{i,j+1}}$. By recursively computing these partial derivatives, we can capture how changes in the current states $D_{i,j}$, $I_{x_{i,j}}$, and $I_{y_{i,j}}$ propagate through the dynamic programming matrix.
C  More results
In the main paper, the main result showcases the performance comparison of LAC and SOTA on the PennAction Dataset [22]. The results represent the average across all 13 actions. Details of these actions can be found in Table 7.
Action Class AP@K Progress $\tau$
10 50 100 K=5 K=10 K=15
baseball_pitch 92.50 90.95 90.68 91.48 91.41 90.95 93.28 88.64
baseball_swing 93.96 91.85 90.36 93.51 93.23 91.8 94.26 95.55
bench_press 94.83 91.64 91.17 93.76 92.88 91.53 96.64 93.42
bowl 96.83 94.53 93.49 92.52 91.96 90.54 86.32 91.84
clean_and_jerk 92.71 88.83 87.07 87.36 86.82 85.64 92.69 94.66
golf_swing 96.37 93.76 93.5 93.28 92.92 91.82 95.86 94.83
jumping_jacks 95.30 94.3 93.87 94.36 95.31 94.66 95.68 96.47
pushup 94.04 94.84 94.38 94.66 94.38 94.23 96.8 96.64
pullup 95.56 94.55 94.26 94.57 94.87 94.12 96.4 95.44
situp 98.52 96.57 96.6 95.22 94.98 94.95 96.25 96.31
squat 97.65 94.28 96.21 95.26 95.52 94.64 93.91 95.08
tennis_forehand 97.51 97.06 97.51 96.55 95.03 94.46 94.29 92.05
tennis_serve 96.62 96.09 95.09 97.76 95.01 95.10 92.33 92.35
average 95.57 93.79 93.40 93.87 93.41 92.65 94.21 94.10
Table 7: Performance comparison of LAC and SOTA on the PennAction Dataset [22] across 13 actions.